OpenStudy (anonymous):

will you please help me with this one find the y intercept and the x intercept for the equation -6x+y=39

OpenStudy (anonymous):

If you want to find the x intercept, set y=0 -6x+0=39 x=-6.5 if you want to find the y intercept, set x=0 -6(0)+y= 39 y=39